OpenStudy (anonymous):

Solve the following problem. It may be helpful to draw a chart on scrap paper to organize the information and write the equation. Be sure to show all steps (V.E.S.T.) and work in order to receive full credit. A clerk was asked to change a $10 bill. She returned 9 more dimes than nickels and twenty-one more quarters than dimes. How many coins of each did she return?

OpenStudy (anonymous):

hmmm that's a puzzler...

OpenStudy (aric200):

1. x = nickels x + 9 = dimes (x + 9) + 21 = quarters 0.05x + 0.10(x + 9) + 0.25(x + 30) = 10.00 5x + 10(x + 9) + 25(x + 30) = 1000 5x + 10x + 90 + 25x + 750 = 1000 40x + 840 = 1000 40x = 160 x = 6 nickels x + 9 = 15 dimes x + 30 = 36 quarters